A plane is going to north with a speed of 150 m/s it is blown by a strong wind at 500 m/s going to west. What is the magnitude of plane's resultant velocity?
1
Expert's answer
2021-09-19T09:25:40-0400


Let's denotne the resultant velocity magnitude as "v". Then from the right triangle shown above have:


"v = \sqrt{v_N^2 + v_W^2} = \sqrt{150^2 +500^2}\approx 522m/s"

Answer. 522 m/s.